# Prompt Public Large Language Models to Synthesize Data for Private On-device Applications

### v1

@@ 1 Introduction
While recent advances of machine learning models significantly benefit from scaling up both training data and model size [24, 2, 32, 15, 39], smaller models have advantages in practical deployment due to inference latency, service cost, and privacy benefits when hosted on the local devices. User data are particularly effective to improve the performance of relatively small models targeting a specific task [18, 47, 7]. Privacy-preserving methods are necessary for training these models using real user data [3]. Differential privacy (DP) [11, 12], a mathematical guarantee applied to characterize the learning process, is a widely acknowledged method to prevent models from memorizing individual user’s information in the training data. Cross-device federated learning (FL) [29, 23], where devices collaboratively learn a model without transferring user data, is popular for limiting data access.
@@ 4 Experiments
As described in [18, 47], our on-device LM is a one-layer LSTM [19, 16] with 670 hidden units, embedding dimension 96, and a 30K-size word-level vocabulary. The LM has about 6M parameters. As shown in Figure 1, we follow [47] to train the LM in two steps: 1) server-side pre-training, and 2) fine-tuning with FL and DP. In DP FL fine-tuning, we run on a production FL system similar to [4] with the real user data. Experiments are performed over two populations of the mobile devices: United States (i.e., the “en-US” population) and India (i.e., the “en-IN” population). To participate in a training round in cross-device FL system, the mobile devices have to satisfy local criteria such as being charging and connecting to unmetered network [4, 21], and minimum separation time across rounds [47]. The total number of available devices are estimated to be around 13M for en-US and 8M for en-IN. Note that the exact population size is unknown as devices are not tracked or logged in the system.
@@ 4.1 Pre-training with Public Data on the Server
Discussion on the quality measure. Evaluating the quality of the generated data is a common challenge. It is even more challenging when the target domain is the private user data, because we need to be extremely careful on privacy protection. In this work, we measure the data quality by training an on-device LM and using federated evaluation to aggregate a single scalar value of NWP accuracy. Developing practical privacy-preserving methods to measure the data quality between the server-side data and the private on-device data is an important future work.
@@ 4.2 Fine-tuning with Differentially Private Federated Learning
After pre-training on the server in Section 4.1, we follow the recommended practices in [47] to fine-tune the LM using a production FL system [4]. For more details about the federated training algorithm and DP accounting, see Appendix B. As the FL training proceeds, we run federated evaluation of the trained LMs over a different set of devices (i.e., the holdout set) in the same population, and report the NWP evaluation metrics in Figure 4. We highlight the following observations:
@@ 5 Improve Synthetic Data with Fine-tuned On-device LM
Note that the goal of this study is to use the fined-tuned LMs to filter the synthetic dataset to get better proxy data for the private domain. We only pre-train on the new proxy data and measure the quality improvement following Section 4.1, but do not further fine-tune the pre-trained model with DP FL as in Section 4.2 and in [42]. Training two models with DP FL on the same population not only complicates the standard production pipeline, but also needs to carefully allocate privacy budgets because the two models are accessing the same set of users (see [42] for more discussions). Although the new proxy data is not used for improving the on-device LMs for Gboard, it can be of independent interest for other tasks. For example, the proxy data can be used to facilitate research simulation in the datacenter, and improve server-side models that cannot be easily fine-tuned with DP FL.
@@
  - Filter C4 17GB (12% LLM-filter-C4-136G)
  - Syn chat 15GB (52% LLM-syn-chat-29G)
@@ 6 Conclusion
Inspired by recent advances in generative LLMs, this work studies whether LLMs can help differentially private federated learning for training small on-device models. In particular, we focus on Gboard (Google Keyboard, a production mobile keyboard application) where the goal is to learn a small on-device LM using the private user typing data. Our work provides evidence that even with no access to the private data, common sense knowledge and careful prompt design can help guide LLMs to synthesize data similar to the target domain. Effectiveness of our method is verified by extensive FL experiments over the real-world user data from the millions of mobile devices.
@@ Appendix B Formal DP Definition and DP FL Algorithm
The Federated Averaging (FedAvg) algorithm and its variants  [29, 43] are widely used in practical FL systems. In training round $i$ of generalized FedAvg, the server first broadcasts a global model $w_{i}$ to a subset of clients; each client $j$ then updates their local model $\theta$ with their local data, typically by a SGD optimizer, and sends back the model delta between learned and initial local model weights $\Delta_{i}^{j}=\theta_{i}^{j}-w_{i}$;the model deltas are aggregated $\Delta_{i}=\sum_{j}\Delta_{j}$ and used as pseudo gradient on the server to update the global model. After training typically thousands of rounds, a final model can be deployed on mobile devices. DP is achieved by clipping the $l_{2}$ norm of the model delta to control the contribution of each client, and then adding noise to the aggregated deltas on the server. The recent DP-Follw The Regularized Leader (DP-FTRL) [22, 8] algorithms add correlated noise across rounds to achieve strong privacy and utility at the same time, which are used in training production on-device LMs in cross-device FL [47]. We use the tree aggregation DP-FTRL [22] in the generalized FedAvg following  [47], as detailed in algorithm 1.
@@
  Initialize DP tree state $\mathcal{T}$ with $z$
@@
     $\bar{P}^{t}=\beta\bar{P}^{t-1}+P^{t}$, $w^{t}\leftarrow w^{0}+\eta_{s}\bar{P}^{t}$   function ClientUpdate($i$, $\theta$      $\mathcal{G}\leftarrow$ (user $i$’s local data split into batches)      for batch $g\in\mathcal{G}$ do         $\theta\leftarrow\theta-\eta_{c}\nabla\ell(\theta;g)$      $\Delta\leftarrow\theta-\theta_{0}$      $\Delta^{\prime}\leftarrow\Delta\cdot\min{\left(1,\frac{C_{\theta}}{||\Delta||}\right)}$      return $\Delta^{\prime}$
@@ C.3 Reporting privacy guarantees
This section clarifies the nuances of the DP guarantees following the guidelines outlined in [35, Sec. 5.3]. The main differences compared to [47] is the usage of the public data in pre-training, and the exact DP guarantees for model training. We include the entire checklist for completeness.
@@ Appendix D More Experimental Results
  - Filter C4 18GB (13% LLM-filter-C4-136G)
  - Syn chat 14GB (48% LLM-syn-chat-29G)

### v2

@@ 1 Introduction
While recent advances of machine learning models significantly benefit from scaling up both training data and model size [26, 2, 35, 15, 42], smaller models have advantages in practical deployment due to inference latency, service cost, and privacy benefits when hosted on the local devices. User data are particularly effective to improve the performance of relatively small models targeting a specific task [18, 51, 7]. Privacy-preserving methods are necessary for training these models using real user data [3, 59]. Differential privacy (DP) [11, 12], a mathematical guarantee applied to characterize the learning process, is a widely acknowledged method to prevent models from memorizing individual user’s information in the training data. Cross-device federated learning (FL) [32, 25], where devices collaboratively learn a model without transferring user data, is popular for limiting data access.
@@ 2 Background
Generating synthetic data to resemble the private domain in the FL setting is an active research area. Several concurrent papers [22, 29] explore data synthesis methods that require more interactions between the server and the private devices. This paper studies a simple prompt engineering approach, and provides unique value by examining the performance of synthesizing data in real-world production FL experiments.
@@ 4 Experiments
As described in [18, 51], our on-device LM is a one-layer LSTM [19, 16] with 670 hidden units, embedding dimension 96, and a 30K-size word-level vocabulary. The LM has about 6M parameters. As shown in Figure 1, we follow [51, 50] to train the LM in two steps: 1) server-side pre-training, and 2) fine-tuning with FL and DP. In DP FL fine-tuning, we run on a production FL system [4] with the real user data. Experiments are performed over two populations of the mobile devices: United States (i.e., the “en-US” population) and India (i.e., the “en-IN” population).
To participate in a training round in cross-device FL system, the mobile devices have to satisfy local criteria such as being charging and connecting to unmetered network [4, 23], and minimum separation time across rounds [51]. The total number of available devices are estimated to be around 13M for en-US and 8M for en-IN. Note that the exact population size is unknown as devices are not tracked or logged in the system.
@@ 4.1 Pre-training with Public Data on the Server
Discussion on the quality measure. Evaluating the quality of the generated data is a common challenge. It is even more challenging when the target domain is the private user data, because we need to be extremely careful on privacy protection. In this work, we measure the data quality by training an on-device LM and using federated evaluation to aggregate a single scalar value of NWP accuracy. Developing practical privacy-preserving methods to measure the distribution similarity between the server-side data and the private on-device data is an important future work. A related work is [21], which proposes a method FreD to measure the distance between a server-side dataset and the private on-device data, and use it to select the best server-side dataset for pre-training. Applying FreD in practice and combining with our synthetic data approach is an interesting direction for future work. Note that this requires careful allocation of privacy budgets in FreD and federated fine-tuning.
@@ 4.2 Fine-tuning with Differentially Private Federated Learning
After pre-training on the server in Section 4.1, we follow the recommended practices in [51] to fine-tune the LM using a production FL system [4]. For more details about the federated training algorithm and DP accounting, see Section 2 and Appendix B.
As the FL training proceeds, we run federated evaluation of the trained LMs over a different set of devices (i.e., the holdout set) in the same population, and report the NWP evaluation metrics in Figure 4. We highlight the following observations:
@@ 5 Improve Synthetic Data with Fine-tuned On-device LM
- Filter C4 17GB (12% LLM-filter-C4-136G)
- Syn chat 15GB (52% LLM-syn-chat-29G)
@@
Note that the goal of this preliminary study is not training new on-device LMs as in Section 4.2, instead, the goal is to see if we can use an existing privately-trained LM to obtain a better server-side dataset. This new dataset can be of independent interest for server-side tasks such as research simulation in the datacenter, and improving server-side models that cannot be easily fine-tuned with DP FL. The process of obtaining this new server-side dataset (i.e., using an existing privately-trained LM to filter the original data) does not incur extra privacy cost, following the post-processing property of differential privacy. We did not use this new dataset to further pre-train an on-device LM, and fine-tune it with DP FL as in Section 4.2. We can potentially adopt the mid-training approach in  [45], which will complicate the standard production pipeline and need specialized privacy accounting.
@@ 6 Conclusion
Inspired by recent advances in generative LLMs, this work studies how LLMs can help differentially private federated learning for training small on-device models. In particular, we focus on Gboard (Google Keyboard, a production mobile keyboard application) where the task is to learn a small on-device LM using the private user typing data. Our work provides evidence that even with no access to the private data, common sense knowledge and careful prompt design can help guide LLMs to synthesize data similar to the target domain. Effectiveness of our method is verified by extensive FL experiments over the real-world user data from the millions of mobile devices.
@@ Appendix B Formal DP Definition and DP FL Algorithm
The Federated Averaging (FedAvg) algorithm and its variants  [32, 46] are widely used in practical FL systems. In training round $t$ of generalized FedAvg, the server first broadcasts a global model $w^{t}$ to a subset of clients; each client $i$ then updates their local model $\theta$ with their local data, typically by a SGD optimizer, and sends back the model delta between learned and initial local model weights $\Delta_{i}^{t}=\theta_{i}^{t}-w^{t}$;the model deltas are aggregated $\Delta^{t}=\sum_{i}\Delta_{i}^{t}$ and used as pseudo gradient on the server to update the global model. After training typically thousands of rounds, a final model can be deployed on mobile devices. DP is achieved by clipping the $l_{2}$ norm of the model delta to control the contribution of each client, and then adding noise to the aggregated deltas on the server. The recent DP-Follw The Regularized Leader (DP-FTRL) [24, 8] algorithms add correlated noise across rounds to achieve strong privacy and utility at the same time, which are used in training production on-device LMs in cross-device FL [51]. We use the tree aggregation DP-FTRL [24] in the generalized FedAvg following  [51], as detailed in algorithm 1.
@@
  Initialize DP tree state $\mathcal{T}$ with $zC$
@@
     $\bar{P}^{t}=\beta\bar{P}^{t-1}+P^{t}$, $w^{t}\leftarrow w^{0}+\eta_{s}\bar{P}^{t}$mdframed}[ linecolor=black, linewidth=1.5pt, roundcorner=4pt, userdefinedwidth=]
  function ClientUpdate($i$, $\theta$
     $\mathcal{G}\leftarrow$ (user $i$’s local data split into batches)
     for batch $g\in\mathcal{G}$ do
        $\theta\leftarrow\theta-\eta_{c}\nabla\ell(\theta;g)$
     $\Delta\leftarrow\theta-\theta_{0}$
     $\Delta^{\prime}\leftarrow\Delta\cdot\min{\left(1,\frac{C}{||\Delta||}\right)}$
     return $\Delta^{\prime}$
@@ C.3 Reporting privacy guarantees
This section clarifies the nuances of the DP guarantees following the guidelines outlined in [38, Sec. 5.3]. The main differences compared to [51] is the usage of the public data in pre-training, and the exact DP guarantees for model training. We include the entire checklist for completeness.
@@ Appendix D More Experimental Results
- Filter C4 18GB (13% LLM-filter-C4-136G)
- Syn chat 14GB (48% LLM-syn-chat-29G)